Two blocks M=1kg and m=0.9kg are connected by a string. The big block M slides on the frictionless table while the block m, hung passing over a pulley by the edge of the table, is falling. What is the acceleration of M?
Answer #1

M=1KG   ,m=0.9kg

frictionless table no need to consider friction

T= Ma

a=(m/M+m)g

a=(0.9/1+0.9)9.8 =4.6m/s2